OpenStudy (anonymous):

Factor the following quadratic 4x^3-x^2-4x+1

OpenStudy (anonymous):

this is cubic not quadratic

OpenStudy (anonymous):

well its on my quadratics quiz so...

OpenStudy (31356):

My homework is to factor things out too lol :D

OpenStudy (anonymous):

I just need help :(

OpenStudy (31356):

Me too :D

OpenStudy (anonymous):

im not sure how to do this correctly

OpenStudy (raden):

4x^3-x^2-4x+1 put x=1, get 4(1)^3-(1)^2-4(1)+1 = 4 - 1 - 4 + 1 = 0 it means (x-1) is one of the factor of 4x^3-x^2-4x+1. then you can find the other factors by using long division or by using Horner

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

(x^2-1)(4x-1), but you can factor further to (4x-1)(x+1)(x-1)

OpenStudy (raden):

yeah, that's full factor for it. but it would be nice if you can show your steps (not only answer) @Nanalew

OpenStudy (anonymous):

im not so sure how you got that but thanks

OpenStudy (anonymous):

Okay then, trying to factor by pulling out so split the expression into two groups: Group 1: -x2+1 Group 2: 4x3-4x Pull out from each group separately : Group 1: (-x2+1) • (1) = (x2-1) • (-1) Group 2: (x2-1) • (4x) Add up: (x^2-1)(4x-1) Then factor (x^2-1) Check : 1 is the square of 1 Check : x2 is the square of x1 Factorization is : (x + 1) • (x - 1) Then (4x - 1) • (x + 1) • (x - 1)

OpenStudy (anonymous):

Heres the work @RadEn, sorry should have posted before answer :)

OpenStudy (raden):

ok, nope. better late than never :)